OpenStudy (anonymous):

Find a point-slope of the line having the given slope and containing the given point. m=3/4, (1,6) can someone help with this question. Thank you.

OpenStudy (anonymous):

Use the equation of a line y=mx+b. You have x,y and the slope (m), you just need to solve for b, so: y = mx +b 6 = (3/4)*1 + b b = 6 - (3/4)*1 b = 21/4 Place b back in the original equation: so the equation becomes y = (3/4)x + 21/4

OpenStudy (anonymous):

Thank you kindly for your help.